# RLHF and IIA: Perverse Incentives

### 02 Dec 2023 v1

@@ A.1 RLPO
Under Assumption 1, for all $K\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{K}}{1-\zeta^{K}-(1-\zeta)^{K}}$, then there exists $\epsilon>0$ such that for all $\delta>0$, there exists a $\tilde{N}>0$ such that for all $N>\tilde{N}$,
($\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)>1-\delta.$
@@
With a little bit abuse of notation, we define
[…]
where $[N]$ is a shorthand notation for $\{1,2,\ldots,N\}$. Note that $\mathcal{I}$ is the set of indices of data that contain both categories of messages. We also define
$\epsilon_{0}$ $=F(P_{\overline{\pi}}(\mathcal{M}_{1}))-p_{*}(1)>0$
$q_{1}^{(n)}$ $=\mathbbm{1}(y^{(n)}\in\mathcal{M}_{1}),$
then by definition $q_{1}=\frac{1}{N}\sum_{n=1}^{N}q_{1}^{(n)}$.
For $n\in[N]$, $k_{1}^{(n)}\sim\mathrm{Binomial}(K,P_{\overline{\pi}}(\mathcal{M}_{1}))$. For $n\in\mathcal{I}$, the conditional mean
$\mathbb{E}\left[k_{1}^{(n)}\mid n\in\mathcal{I}\right]=\frac{KP_{\overline{\pi}}(\mathcal{M}_{1})-KP_{\overline{\pi}}(\mathcal{M}_{1})^{K}}{1-P_{\overline{\pi}}(\mathcal{M}_{1})^{K}-P_{\overline{\pi}}(\mathcal{M}_{2})^{K}}=K\cdot F(P_{\overline{\pi}}(\mathcal{M}_{1})).$
On the other hand, the distribution of $q_{1}^{(n)}$ satisfies
$q_{1}^{(n)}=\begin{cases}1&\text{if }k_{1}^{(n)}=K\\ 0&\text{if }k_{1}^{(n)}=0\\ \mathrm{Bernoulli}(p_{*}(1))&\text{otherwise},\end{cases}$
so its conditional mean $\mathbb{E}\left[q_{1}^{(n)}\mid n\in\mathcal{I}\right]=p_{*}(1)$. We can decompose $q_{1}$ into two sums as
$q_{1}=\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}+\frac{1}{N}\sum_{n\notin\mathcal{I}}q_{1}^{(n)}=\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}+\frac{1}{NK}\sum_{n\notin\mathcal{I}}k_{1}^{(n)}.$ (21)
For any fixed subset $I\subset[N]$, by Hoeffding’s inequality, we have
$\mathbb{P}\left(\frac{1}{|I|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-F(P_{\overline{\pi}}(\mathcal{M}_{1}))\leq-\frac{\epsilon_{0}}{2}\right)\leq e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}.$ (22)
Similarly,
$\mathbb{P}\left(\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}-p_{*}(1)\geq\frac{\epsilon_{0}}{4}\right)$$\leq$ $e^{-\frac{|I|\epsilon_{0}^{2}}{8}}$ (23)
$\mathbb{P}\left(\frac{1}{|I|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}\leq\frac{p_{*}(1)}{2}\right)$$\leq$ $e^{-\frac{|I|p_{*}(1)^{2}}{2}}.$ (24)
@@
($\quad\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}-\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>F(P_{\overline{\pi}}(\mathcal{M}_{1}))-\frac{\epsilon_{0}}{2}>p_{*}(1)+\frac{\epsilon_{0}}{4}>\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>\frac{\epsilon_{0}}{2}+p_{*}(1)>p_{*}(1)+\frac{\epsilon_{0}}{4}>\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}>\frac{\epsilon_{0}}{2}+p_{*}(1)\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}<p_{*}(1)+\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{p_{*}(1)}{2}\right\}\right)$
$>1-e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{|I|\epsilon_{0}^{2}}{8}}-e^{-\frac{|I|p_{*}(1)^{2}}{2}},$ (25)
where in the first inequality we used the fact that $\epsilon_{0}=F(P_{\bar{\pi}}(\mathcal{M}_{1}))-p_{*}(1)$. By Lemma 8, there exists $10<\gamma<$ 1 such that $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$. By the law of total probability,
@@
($\quad\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)$
$=\sum_{I\in 2^{[N]}}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\bigcap\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\mid\left\{\mathcal{I}=I\right\}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$=\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\mathbb{P}\left(\left\{\frac{1}{|I|K}\sum_{n\in I}k_{1}^{(n)}-\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|I|}\sum_{n\in I}q_{1}^{(n)}>0\right\}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right).$
Applying the union bound in (25),
@@
($\quad\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)$
$\geq\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\left(1-e^{-\frac{|I|K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{|I|\epsilon_{0}^{2}}{8}}-e^{-\frac{|I|p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$\geq\sum_{I\in 2^{[N]}:|I|>N(1-\gamma)}\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(\left\{\mathcal{I}=I\right\}\right)$
$=\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\mathbb{P}\left(|\mathcal{I}|>N(1-\gamma)\right)$
$\geq\left(1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}\right)\cdot\left(1-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}\right)$
$\geq 1-e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}-e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}-e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}.$
We let $\delta>4e^{-\frac{N(1-\gamma)\gamma^{2}\epsilon_{0}^{2}p_{*}(1)^{2}}{8}}>e^{-\frac{N(1-\gamma)K^{2}\epsilon_{0}^{2}}{2}}+e^{-\frac{N(1-\gamma)\epsilon_{0}^{2}}{8}}+e^{-\frac{N(1-\gamma)p_{*}(1)^{2}}{2}}+e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$, which in turn gives $N>\tilde{N}=\left\lceil\frac{8\ln(4/\delta)}{(1-\gamma)\gamma^{2}\epsilon_{0}^{2}p*(1)^{2}}\right\rceil$. Let $\epsilon=\frac{(1-\gamma)\epsilon_{0}}{4}$. Then for all $N>\tilde{N}$,
($\quad\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{N}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{(1-\gamma)\epsilon_{0}}{4}\right\}\bigcap\left\{q_{1}>0\right\}\right)$
=($=\mathbb{P}\left(\left\{\frac{|\mathcal{I}|}{N}\left[\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}\right]>\frac{(1-\gamma)\epsilon_{0}}{4}\right\}\bigcap\left\{q_{1}>0\right\}\right)$
@@
($\geq\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|K}\sum_{n\in\mathcal{I}}k_{1}^{(n)}-\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>\frac{\epsilon_{0}}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}q_{1}^{(n)}>0\right\}\bigcap\left\{|\mathcal{I}|>N(1-\gamma)\right\}\right)>1-\delta.$
Step 3: RLPO produces higher rewards for category 2 messages with high probability.
The last step combines results from the previous two steps, showing that under the stated assumptions, the reward model obtained via RLPO assigns a higher reward to the category of messages preferred by the minority of the population with high probability.
Lemma 10.
Under Assumptions 1 and 2, for all $K\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{K}}{1-\zeta^{K}-(1-\zeta)^{K}}$, and $\widehat{\psi}$ is as defined in Proposition 4, then there exists $\eta>0$ such that
$\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1.$
@@
By Lemma 9, there exists $\epsilon>0$ such that for all $\delta>0$, there exists a $\tilde{N}>0$ such that for all $N>\tilde{N}$,
($\mathbb{P}\left(\left\{\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}>\epsilon\right\}\bigcap\left\{q_{1}>0\right\}\right)>1-\delta.$
Note that, under the high probability event, $\widehat{\psi}=\operatorname*{arg\,min}_{\psi}\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$. By Lemma 7, with probability $1-\delta$,
$\psi_{2}-\psi_{1}>1-\left[1+K^{2}\cdot\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right)\right]^{-1}>1-\frac{1}{1+\epsilon}=\frac{\epsilon}{1+\epsilon}.$
Here, the denominator $\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}\neq 0$ under this high probability event. Setting $\eta=\frac{\epsilon}{1+\epsilon}>0$, with probability $1-\delta$, $\psi_{2}-\psi_{1}>\eta$ for all $N>\tilde{N}$. In other words, for all $K\geq 3$, there exists $\eta>0$ such that for all $\delta>0$, $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)\geq 1-\delta$. Since $\eta$ does not depend on $\delta$, taking $\delta\to 0$ gives $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1$.
Finally, we prove the following failure case for RLPO.
See 4
Proof.
By Lemma 10, there exists $\eta>0$ such that $\lim_{N\to\infty}\mathbb{P}(\psi_{2}-\psi_{1}>\eta)=1$ with probability $1-\delta$.
When $\beta=0$, $\widehat{\pi}_{\theta}$ maximizes reward. Lemma 10 equivalently states that for all $\delta>0$, there exists $\tilde{N}>0$ such that for all $N=|\mathcal{D}|>\tilde{N}$, with probability $1-\delta^{\prime}$, $\psi_{2}-\psi_{1}>\eta$. This implies that $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})=0$ and $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})=1$. Equivalently, for all $\delta>0$, $\lim_{N\to\infty}P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})=0$ with probability $1-\delta$. Since this holds for arbitrary $\delta$, it follows that
$P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})\stackrel{{\scriptstyle p.}}{{\longrightarrow}}0\qquad\text{and}\qquad P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})\stackrel{{\scriptstyle p.}}{{\longrightarrow}}1.$
When $\beta>0$, $\widehat{\pi}_{\theta}$ satisfies
$P_{\widehat{\pi}_{\theta}}(x)=\frac{P_{\overline{\pi}}(x)e^{\widehat{r}(x)N/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{\widehat{r}(x^{\prime})N/\beta}}.$ (26)
Summing over all messages of category 1 and 2 respectively,
$\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}=\frac{\sum_{x\in\mathcal{M}_{1}}P_{\widehat{\pi}_{\theta}}(x)}{\sum_{x^{\prime}\in\mathcal{M}_{2}}P_{\widehat{\pi}_{\theta}}(x^{\prime})}=\frac{\sum_{x\in\mathcal{M}_{1}}P_{\overline{\pi}}(x)}{\sum_{x^{\prime}\in\mathcal{M}_{2}}P_{\overline{\pi}}(x^{\prime})}\cdot e^{\left(\psi_{1}-\psi_{2}\right)N/\beta}=\frac{p_{1}}{p_{2}}\cdot e^{\left(\psi_{1}-\psi_{2}\right)N/\beta}.$
Lemma 10 implies that, for all $\delta>0$, there exists $\tilde{N}>0$ such that for all $N>\tilde{N}$, with probability $1-\delta$, $\psi_{2}-\psi_{1}>\eta$. Thus,
$\psi_{1}-\psi_{2}$ $<-\eta\implies\left(\psi_{1}-\psi_{2}\right)\cdot N/\beta<-\frac{\eta N}{\beta}.$ (27)
Therefore, the policy ratio can be bounded by
$\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}.$
Thus, for all $\delta>0$,
$\lim_{N\to\infty}\mathbb{P}\left(\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}\right)\geq 1-\delta.$
Thus, $\lim_{N\to\infty}\mathbb{P}\left(\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}<\frac{p_{1}}{p_{2}}\cdot e^{-\frac{\eta N}{\beta}}\right)=1$, which implies that $\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}\overset{p.}{\longrightarrow}0$, and the proof follows by noticing that $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})+P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})=1$. ∎

### 21 Dec 2023 v2

@@ A.1 RLPO
Under Assumption 1, for all $|\mathcal{Y}|\geq 3$, if $p_{*}(1)<F(P_{\overline{\pi}}(\mathcal{M}_{1}))$ with $F(\zeta)=\frac{\zeta-\zeta^{|\mathcal{Y}|}}{1-\zeta^{|\mathcal{Y}|}-(1-\zeta)^{|\mathcal{Y}|}}$, then there exists $\eta>0$ such that $\mathbb{P}(\mathcal{E}_{\eta}(\mathcal{D}))\to 1$ as $|\mathcal{D}|\to\infty$.
@@
Let $\delta=F(P_{\overline{\pi}}(\mathcal{M}_{1}))-p_{*}(1)>0$. Under Assumption 1, $|\mathcal{Y}\cap\mathcal{M}_{1}|\sim\mathrm{Binomial}(|\mathcal{Y}|,P_{\overline{\pi}}(\mathcal{M}_{1}))$, giving
$\mathbb{E}\left[\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}\mid(\mathcal{Y},y)\in\mathcal{I}\right]=\frac{P_{\overline{\pi}}(\mathcal{M}_{1})-P_{\overline{\pi}}(\mathcal{M}_{1})^{|\mathcal{Y}|}}{1-P_{\overline{\pi}}(\mathcal{M}_{1})^{|\mathcal{Y}|}-P_{\overline{\pi}}(\mathcal{M}_{2})^{|\mathcal{Y}|}}=F(P_{\overline{\pi}}(\mathcal{M}_{1})).$
On the other hand,
$\mathbb{E}\left[\mathbbm{1}(y\in\mathcal{M}_{1})\mid(\mathcal{Y},y)\in\mathcal{I}\right]=p_{*}(1).$
For $(\mathcal{Y},y)\notin\mathcal{I}$, observe that $\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}=\mathbbm{1}(y\in\mathcal{M}_{1})$. Thus
$\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}-\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})=\frac{|\mathcal{D}|}{|\mathcal{I}|}(\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}).$ (21)
By Hoeffding’s inequality,
$\mathbb{P}\left(\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}-F(P_{\overline{\pi}}(\mathcal{M}_{1}))\leq-\frac{\delta}{2}\mid\mathcal{I}\right)$ $\leq e^{-\frac{|\mathcal{I}|\delta^{2}}{4}}$
$\mathbb{P}\left(\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})-p_{*}(1)\geq\frac{\delta}{4}\mid\mathcal{I}\right)$ $\leq e^{-\frac{|\mathcal{I}|\delta^{2}}{8}}$
$\mathbb{P}\left(\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})-p_{*}(1)\leq-\frac{p_{*}(1)}{2}\mid\mathcal{I}\right)$ $\leq e^{-\frac{|\mathcal{I}|p_{*}^{2}(1)}{2}}.$
@@
($\quad\mathbb{P}\left(\left\{\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}-\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})>\frac{\delta}{4}\right\}\bigcap\left\{\frac{1}{|\mathcal{I}|}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})>0\right\}\mid\mathcal{I}\right)$
$>1-e^{-\frac{|\mathcal{I}|\delta^{2}}{4}}-e^{-\frac{|\mathcal{I}|\delta^{2}}{8}}-e^{-\frac{|\mathcal{I}|p_{*}^{2}(1)}{2}}.$
By Lemma 10, there exists $10<\gamma<$ 1 such that $\mathbb{P}(|\mathcal{I}|\leq|\mathcal{D}|(1-\gamma))\leq e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}$. By Equation (21),
@@
($\quad\mathbb{P}\left(\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}(\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}})>\frac{\delta}{4}\right\}\bigcap\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}\rho_{\mathrm{chosen}}>0\right\}\bigcap\left\{|\mathcal{I}|>|\mathcal{D}|(1-\gamma))\right\}\right)$
@@
=($=\mathbb{P}\left(\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}(\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}})>\frac{\delta}{4}\right\}\bigcap\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}\rho_{\mathrm{chosen}}>0\right\}\mid|\mathcal{I}|>|\mathcal{D}|(1-\gamma))\right)\mathbb{P}\left(|\mathcal{I}|>|\mathcal{D}|(1-\gamma))\right)$
$>\left(1-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{4}}-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{8}}-e^{-\frac{|\mathcal{D}|(1-\gamma)p_{*}^{2}(1)}{2}}\right)\left(1-e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}\right)$
$>1-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{4}}-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{8}}-e^{-\frac{|\mathcal{D}|(1-\gamma)p_{*}^{2}(1)}{2}}-e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}.$
Let $\eta=\frac{(1-\gamma)\delta}{4}>0$. It follows that
$\mathbb{P}\left(\mathcal{E}_{\eta}(\mathcal{D})\right)$ =($=\mathbb{P}\left(\left\{\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}>\eta\right\}\bigcap\left\{\rho_{\mathrm{chosen}}>0\right\}\right)$
($\geq\mathbb{P}\left(\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}(\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}})>\frac{\delta}{4}\right\}\bigcap\left\{\frac{|\mathcal{D}|}{|\mathcal{I}|}\rho_{\mathrm{chosen}}>0\right\}\bigcap\left\{|\mathcal{I}|>|\mathcal{D}|(1-\gamma))\right\}\right)$
$>1-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{4}}-e^{-\frac{|\mathcal{D}|(1-\gamma)\delta^{2}}{8}}-e^{-\frac{|\mathcal{D}|(1-\gamma)p_{*}^{2}(1)}{2}}-e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}.$
The proof follows by taking $|\mathcal{D}|\to\infty$. ∎
We are now ready to prove the following failure case for RLPO. See 4
@@
The proof follows from applying Lemmas 9 and 11. ∎